# Heart-felt Narratives: Tracing Empathy and Narrative Style in Personal Stories with LLMs

### v1

@@ 9 Conclusion
In this work, we define and quantify narrative style as it relates to narrative empathy. We introduce Heart, the first theory-driven taxonomy delineating elements of narrative style that can evoke empathy towards a story. We evaluate the performance of LLMs in extracting narrative elements from Heart, showing that prompting with our taxonomy leads to reasonable, human-level annotations beyond what prior lexicon-based methods can do, but that LLMs struggle in specific tasks, such as extracting expressions of cognition and evaluations. Through a crowdsourced study with over 2,000 participants, we demonstrate how Heart can be used to empirically understand the empathic role of narrative style factors. We find that vividness of emotions expressed, character development and plot volume are related to narrative empathy, and contextual factors such as a person’s baseline trait empathy or sharing an experience with the narrator contribute to these effects. Additionally, we show that empathy responses are highly variable even in the same story, and that narrative style preferences are personalized to people with different demographic profiles (such as varying levels of trait empathy). Our findings show the promise of using LLMs for annotating complex story features that can yield interesting social and behavioral insights.
@@
Personal stories can contain intimate and vulnerable information, in addition to inducing emotions in readers. Our study protocol for showing sensitive stories to crowdworkers was approved by our institution’s ethics review board as an exempt study. We ensured that all datasets we used were also collected via IRB-approved protocols.
@@ Appendix B Distribution of Empathy Standard Deviation
Appendix C Codebook and GPT-4 Prompts
@@ C.10 Vividness of setting
Appendix D Error Analysis Story Examples
@@ D.2 Stories with Cognition Disagreement
Human: 0.25 GPT: 1.0 I’ve been hearing a lot of people saying that MIT students aren’t successful as Stanford or Harvard students because there aren’t as many well-known MIT CEOs. It seems rather unfair of them to say that because MIT students have contributed a lot to this world from nobel-prize winning theorems to groundbreaking algorithms.

### v2

@@ 9 Conclusion
In this work, we quantify narrative style as it relates to narrative empathy. We introduce Heart, the first theory-driven taxonomy delineating elements of narrative style that can evoke empathy towards a story. We evaluate the performance of LLMs in extracting narrative elements from Heart, showing that prompting GPT-4 with our taxonomy leads to reasonable, human-level annotations beyond what prior lexicon-based methods can do, but that LLMs struggle in specific tasks, such as GPT-4’s limited ability to extract expressions of cognition and evaluations. Through a crowdsourced study with over 2,000 participants, we demonstrate how Heart can be used to empirically understand the empathic role of narrative style factors. We find that vividness of emotions expressed, character development and plot volume are related to narrative empathy, and contextual factors such as a person’s baseline trait empathy or sharing an experience with the narrator contribute to these effects. Additionally, we show that empathy responses are highly variable even in the same story, and that narrative style preferences are personalized to people with different demographic profiles (such as varying levels of trait empathy). Our findings show the promise of using LLMs for annotating complex story features that can yield interesting social and behavioral insights.
@@
Personal stories can contain intimate and vulnerable information, in addition to inducing emotions in readers. Our study protocol for showing sensitive stories to crowdworkers was approved by our institution’s ethics review board as an exempt study. Participants gave informed consent that their survey ratings would be collected via Prolific. We ensured that all datasets we used were also collected via IRB-approved protocols, and will only distribute our dataset to IRB-approved protocols.
@@ Appendix B Distribution of Empathy Standard Deviation
Appendix C Codebook and LLM Prompts
@@ C.10 Vividness of setting
Appendix D GPT-4 Error Analysis Story Examples
Scores range from 0 to 1, where 0 indicates low presence of narrative feature and 1 indicates high presence.
@@ D.2 Stories with Cognition Disagreement
Human: 0.25
GPT-4: 1.0
I’ve been hearing a lot of people saying that MIT students aren’t successful as Stanford or Harvard students because there aren’t as many well-known MIT CEOs. It seems rather unfair of them to say that because MIT students have contributed a lot to this world from nobel-prize winning theorems to groundbreaking algorithms.